OpenStudy (anonymous):

Use the substitution method to solve the system of equations. 2x – 5y = –14 x – 2y = –5

OpenStudy (anonymous):

x-2y+-5 x=2y-5 <-- plug that in to the other equation 2(2y-5)-5y=-14 then distribute to get, 4y-10-5y=-14 -y-10=-14, then add 10 to both sides -y=-4, then divde by -1 y=4 plug into other equation x=2(4)-5 x=8-3 x=5 so y=4 and x=5